OpenStudy (anonymous):

WILL MEDAL AND FAN ALL WHO ARE RIGHT What is the solution to the equation? 1/3x + 1/3(2x - 15) = 31/2

OpenStudy (anonymous):

@ankit042

OpenStudy (anonymous):

@sangya21 help!

OpenStudy (anonymous):

last question

OpenStudy (ankit042):

(x-3)/x(2x-15) = 31/2 2(x-3) = 31(2x^2-15x)

OpenStudy (anonymous):

1/3x + 1/3(2x - 15) = 31/2 1/3x + 2/3x - 15/3 = 31/2 3/3x - 5 = 31/2 x = 31/2 + 5